OpenStudy (anonymous):

True or False: If synthetic division is used to calculate the quotient of two polynomials, then the degree of the quotient will always be one more than the degree of the dividend.

OpenStudy (anonymous):

@mathmale can you help me?

OpenStudy (mathmale):

What do you need to know here? Are you familiar with the vocabulary, such as "degree," "quotient," "dividend?"

OpenStudy (anonymous):

Yes but I don't really remember everything. I want to know if that's true or false

OpenStudy (mathmale):

I'd like to help, but just giving you the answer will not help you in the long run. Which of the vocabulary is clear and which is not? Please look up the voc. that is unclear. Then we can proceed further.

OpenStudy (anonymous):

Okay. I did

OpenStudy (mathmale):

what do you still need to know?

OpenStudy (anonymous):

well I just don't get the question

OpenStudy (mathmale):

What is the degree of y = x^5 - 4x^3 + 7x -14? of y = x^3 +8?

OpenStudy (anonymous):

-2?

OpenStudy (mathmale):

Happy to help you, so long as you are contributing, looking up vocabulary wherever necessary, asking questions. You have to be involved. "Degree" is not and cannot be negative. In any case, neither of the polynomials I gave you as examples has degree 2. Again, if you do not know the meaning of any of the vocab., please look it up. I have 43 years of experience teaching math and can usually tell when a student is just trying to guess.

OpenStudy (mathmale):

What is the degree of y = x^5 - 4x^3 + 7x -14? of y = x^3 +8? Define "degree" as it applies to this situatino (algebra).

OpenStudy (anonymous):

ohh so the degree is 5 and 3?

OpenStudy (mathmale):

Yes. That's better. Now suppose you divide the 5th degree poly by the 3rd degree poly. What degree will the quotient have? How did you get your answer?

OpenStudy (anonymous):

um I don't remember how to dived degrees

OpenStudy (anonymous):

would it be x^2?

OpenStudy (mathmale):

Define "degree" as it applies to this situation (algebra).

OpenStudy (anonymous):

For a term with one variable, the degree is the variable's exponent.

OpenStudy (mathmale):

Re: "would it be x^2?" No. "Degree" is just positive, integer number; one does not include "x^2" in specifying degree. If you have a textbook, look for an example of synthetic division so that we could discuss something concrete.

OpenStudy (mathmale):

Which is the polynomial with degree 5: the divisor, the dividend, or the quotient? Which is the poly with degree 3: the divisor, the dividend, or the quotient?

OpenStudy (anonymous):

The polynomial with degree 5 is the dividend

OpenStudy (mathmale):

Right. You are dividing that polynomial by the other polynomial. How do you determine the degree the quotient will have? What is that degree?

OpenStudy (anonymous):

Will it be the biggest degree? So degree 5